# Zebrafix: Mitigating Memory-Centric Side-Channel Leakage via Interleaving

### v2

@@ III-B Interleaving as Leakage Defense
Figure 4: Interleaving data and counter values in memory. In order to ensure freshness in each 16-byte block in memory, the blocks are split into 8-byte counter and 8 byte available for data. On each memory write, the counter value is incremented by one.
@@ III-C Realization Considerations
Cipherfix takes the first approach: The authors modify the application source code to call a special dummy function, passing each secret as a parameter. They then employ dynamic taint tracking to find all instructions accessing that variable. Using dynamic methods for secrecy tracking has the main drawback that it is very susceptible to insufficient coverage. Missing a code path during analysis eventually leads to bugs in the protected application, where an instruction accesses a protected memory location under a wrong assumption. Static taint analysis tools may avoid this issue. Yet, static analyses often need to overapproximate results and can still be incomplete in the tracking of exact data types, leading to error-proneness in the instrumentation. Any secrecy tracking that leads to hybrid (instrumented and non-instrumented) usage of global variables endangers functional correctness of the program, as detailed in Section VIII.
@@ IV Defense Implementation Layers
As previous research indicated, ciphertext side-channel defenses can be realized at different implementation layers: The ad-hoc patches to WolfSSL [42] and the Linux kernel [21] were made to the source code, while Cipherfix [37] directly modified the compiled binaries through a binary rewriting framework. In the following, we provide guidance on a suitable implementation layer for an automated mitigation. We accumulate knowledge gained from existing side-channel defenses, explore possible solutions and discuss practical limitations. For compiler-based approaches, we generally use terminology from the LLVM project [19], which has long become a standard tool for side-channel defense implementations. However, our conclusions also apply to similar compiler infrastructures.
@@ IV-A Binary Rewriting
In binary rewriting, the modifications are directly applied on an existing binary. This approach has the advantage of being able to harden an application specifically for deployment on a vulnerable system, e.g., an SEV-SNP protected VM, without needing recompilation. Another advantage is the ability to protect not only a single application binary, but its dependencies as well, which may be too complex to compile and deploy separately (e.g., libc). The most significant challenges of binary rewriting are the stability of the modified binary, and runtime performance [37]. While binary rewriting is generally possible for typical instrumentation tasks, ciphertext side-channel defenses require very extensive changes to the original binary, pushing any binary rewriting framework to its limits [13, 8].
@@ IV-C Compiler: LLVM IR/Middle End
Summary. In order to benefit from eased adjustments to memory layouts, instrumentation needs to be on middle end level. The advantages of back end level instrumentation only become apparent if the instrumentation happens after the register allocation to be able to protect all memory writes. Yet, at the back end stage, it is already sufficiently complicated to adjust the memory layout for interleaving, making it comparable to binary instrumentation. We conclude that hardening ciphertext side-channels at the IR layer promises the best final performance; and combined with a binary-level verification, an appropriate security assurance is provided.
@@ V Zebrafix Implementation
As a case study for evaluating the described interleaving approach, we implement a proof-of-concept, dubbed Zebrafix. With that, we showcase the versatility of interleaving as a side-channel defense. Zebrafix is implemented as a set of LLVM 16111LLVM 16 is at the time of writing the newest LLVM version supported by SVF. out-of-tree passes with SVF [30] pointer analysis, with $2,600$ LOC in total. An overview of the Zebrafix toolchain can be found in Figure 5. The developer who wants to protect a certain function against ciphertext side-channel leakage marks it with a $\mathtt{clang}$ function attribute. The attribute is automatically propagated through the call tree, leading to instrumentation of all called child functions as well. Indirect calls and unsupported components like inline assembly are printed for further inspection by the developer during the call tree instrumentation which is based on the LLVM call graph. We merge all bitcode files that are generated for whole-program analysis and then instrument the merged bitcode file. A discussion about the interaction of instrumented and non-instrumented (e.g., dynamically linked) code can be found in Section VIII. Zebrafix supports the block size 16 byte with data chunks up to size $16/2=8$ byte. An example of data chunks smaller than 8 byte can be found in Figure 6.
@@ V-A Compiler-Level Transformations
In addition to the LLVM IR passes, we adjust the register allocator for x86-64 targets to decrease the number of stack spills of potentially secret data during function calls or because of high register pressure inside of functions. For that, we integrate a register allocator enhancement that spills to vector registers instead of the stack, proposed by Matthias Braun [^3] We adapted https://discourse.llvm.org/t/rfc-spill2reg-selectively-replace-spills-to-stack-with-spills-to-vector-registers/59630/15.. If there are not enough vector registers available for saving values, we print a warning so that the developer can check whether the write actually contains secret-dependent data or just irrelevant data like a pointer that is written to the stack. The check is further enhanced with a binary-level verification tool. For the sake of maintainability and extensibility of Zebrafix, we chose not to implement interleaved stack spilling, which would have required deep changes to the stack allocation.
@@ V-B Adjusting Compiler Intrinsics
Examples of memory-altering intrinsics are $\mathtt{memcpy}$ and $\mathtt{memset}$, which we replace with own implementations in Zebrafix. As the 16-byte blocks that are built with Zebrafix contain varying amounts of data (e.g., 1 byte for chars, 4 byte for short integers, 8 byte for long integers), it does not suffice to adjust the amount of bytes copied or set by a fixed factor. Instead, the correct number of elements of a specific primitive type to be copied has to be determined. In the Zebrafix implementation, we use SVF to infer the types of (source and) destination objects for the memory-altering intrinsics. With that, the data can properly be extracted such that a suitable interleaved struct for the copy or set destination can be generated. The struct is then used for inserting new instrumented writes to memory. In addition, the counter values must not only be copied, but also incremented to prevent leakage from multiple write accesses with the same counter value. Thereby, even with multiple copies of the buffer, the counters are always incremented to avoid repeating ciphertexts.
@@ V-C Instrumentation Aspects for Interleaving
Figure 7: Instrumentation of a 1-byte store instruction. The value to be written is combined with a counter via register-only operations. Finally, a single store instruction writes both value and counter.
@@
V-D Binary-Level Verification
To ensure that the LLVM middle end level instrumentation is not removed by compiler passes that occur later in the compilation process, and that inserted stack spills do not leak any patterns of repeating values, we implement a binary-level verification tool. The verification tool is a Pintool [23] that traces all memory writes of the instrumented binary and manages a set of all written values per address. Whenever there is a repetition of written values within a set, the corresponding instruction offsets and written values are logged to a trace file so that the developer can manually check them for potential leakages in accordance with known leaking algorithms. That way, the developer can determine whether secret data would be spilled to the stack and can apply some source-level mitigations like storing the affected data as global variables instead. Afterward, the compilation and binary-level check is repeated until a non-leaky version is produced (cf. Figure 5).
@@ VI Zebrafix Evaluation
In order to categorize the presented approach of Zebrafix as a mitigation against ciphertext side-channel leakage, we assess the security and performance in comparison to the Cipherfix masking approach. We perform our evaluation on a system with an Intel 4th Gen Xeon Gold 6438Y+ CPU (Sapphire Rapids) with $512$ GiB of DDR5 and local SSD storage. We base our evaluation on target primitives from $\mathtt{libsodium}$-1.0.20 and $\mathtt{mbedTLS}$-3.6.0.
@@ VI-A Compile Time and Memory Consumption
The average code size of instrumented binaries increases by around $63$% (from $220$/$250$ MB to approx. $290$/$435$ MB), where a recurring part is counter management and the rest belongs to other adjustments of instructions and global variables. The average memory usage increases by a factor of five; this is mostly due to lack of optimizations like checking the heap data types or filling each 8-byte data chunk of the interleaved structs with data. The comparatively high memory usage increase in $\mathtt{SHA512}$ is an example of such missing heap data type optimizations. More details can be found in Table I.
Table I: Overhead in terms of compilation time, code size and memory consumption. Compilation time and code size are related to the library version containing certain instrumented primitives (baseline given in the library rows), whereas memory usage is estimated for the execution of the primitive in isolation.
@@
$\mathtt{libsodium}$ $3.4$ $250$
$\mathtt{Ed25519}$ $8.1$ $437$ $2,744$ $10,336$
$\mathtt{SHA512}$ $8.3$ $45$,60056005,6005 , 600 $59,568$
$\mathtt{ChaCha20}$ $7.6$ $389$ $2,288$ $7,488$
$\mathtt{mbedTLS}$ $2.6$ $218$
$\mathtt{Base64}$ $5.8$ $286$ $1,544$ $3,488$
@@ VI-B Runtime Overhead and Security
The primitives for the evaluation were chosen according to the following criteria: $\mathtt{EdDSA}$ has been shown to be vulnerable against ciphertext side-channel attacks [21, 37]. Apart from that, we include $\mathtt{Base64}$ as it was the fastest target in prior ciphertext side-channel mitigation evaluations, and $\mathtt{SHA512}$ as another fast target for comparison. We depict the $\mathtt{Base64}$ decoding in Table II; the results for the encoding are similar. All targets include zeroing of the resp. memory when freeing resources to avoid leakage of data through memory reuse. As mentioned in Section VIII, we exclude target versions that contain inline assembly or features that the current proof-of-concept does not support, like $\mathtt{AES}$ vector instructions. In the latter case, the computation mainly happens in registers, so leaking spots are rare and can be subject to manual adjustments.
@@ VI-C Comparison to Cipherfix Masking Approach
Zebrafix as well as Cipherfix introduce memory overhead from the ciphertext side-channel protection compared to baseline execution. While Zebrafix extends each data chunk with a counter in interleaved memory writes, Cipherfix is based on masking each plaintext chunk via XORing the data and the random mask. In Cipherfix, the Base and Enhanced versions contain a same-sized secrecy and mask buffer for each data chunk. In the Cipherfix-Fast version, there is only one merged buffer for each data chunk. Thereby, the memory overhead in Zebrafix is equal to Cipherfix or even smaller if 8-byte chunks are interleaved, whereas Zebrafix introduces a larger memory overhead when dealing with smaller data types (Figure 6).
@@
Even for the fastest primitive with the weakest security level in the Cipherfix evaluation, $\mathtt{Base64}$ from $\mathtt{mbedTLS}$ and Cipherfix-Fast with $\mathtt{AES}$, Zebrafix reaches a similar overhead. In Cipherfix, only secret-dependent instructions are protected. Thus, Zebrafix’s performance impact could be reduced even further with secret tracking and other optimizations, e.g., allowing the use of vector instructions in combination with interleaved memory writes.
In general, our results suggest that our proposed interleaving approach is more efficient than the data masking approach taken by Cipherfix. This is due to the fact that randomness generation introduces a significant overhead that we avoid. The remaining speedup of Zebrafix over Cipherfix can be explained with a more optimized compiler-built binary, instead of relying on binary rewriting. From this, we conclude that implementing an interleaving-based ciphertext side-channel mitigation in the compiler middle end can enhance performance and security compared to masking-based binary-level approaches.
@@
$\mathtt{Ed25519}$ $39.1$$\times$ $5.5$$\times$ $8.6$$\times$ $1.6$$\times$
$\mathtt{SHA512}$ $21.6$$\times$ $1.6$$\times$ $2.4$$\times$ $1.3$$\times$
@@
$\mathtt{Base64}$ $4.6$$\times$ $1.2$$\times$ $1.6$$\times$ $1.1$$\times$
@@ VII Preventing Leakages from Silent Stores
With the discussion of ciphertext side-channel defenses in mind, we now explore connections to silent stores to show how these memory-centric leakages can be generalized to more holistic defenses. Per se, all attacks described in Section II are based on different mechanisms to exploit leakage. However, despite originating from different sources, all of these leakages can be considered memory-centric side-channels and thus be thwarted by defenses protecting against the whole class: If the data was not architecturally written to memory in a leaky way in the first place, the different leakage mechanisms are no longer effective. In this section, we outline the interconnection of ciphertext side-channel and silent store leakages and compare the effectiveness of Zebrafix with silent store mitigations.
@@
In the ciphertext side-channel leakage scenario, an attacker can observe deterministic patterns of ciphertext changes of victim VM data and thereby derive information about secret keys. The attacker model of ciphertext side-channel leakage is the strongest one to be considered. In the other memory-centric leakage scenarios, we assume that microarchitectural optimizations that are currently on the rise, like the ones discussed in Section II, will gain more importance and see implementation in common processors. We therefore assume a strong attacker to err on the side of caution. The attacker is able to bring the system into a state where the best conditions for maximal leakage through microarchitectural optimizations have been established; thus we assume the worst case leakage in each channel to be possible. [^4] At the time of writing, we are not aware of systems that already feature the microarchitectural optimizations in question, apart from the ones that are linked in Section II. The assumptions on code that is instrumented are in line with the ones presented in Section III-A.

### v3

@@ III-B Interleaving as Leakage Defense
Figure 4: Interleaving data and counter values in memory. In order to ensure freshness in each 16-byte block in memory, the blocks are split into 8-byte counter and 8 byte available for data. On each memory write, the counter value is incremented.
@@ III-C Realization Considerations
Cipherfix takes the first approach: The authors modify the application source code to call a special function, passing each secret as parameters. They then employ dynamic taint tracking to find all instructions accessing that variable. However, using dynamic methods for secrecy tracking is susceptible to insufficient coverage. Missing a code path during analysis eventually leads to bugs in the protected application, where an instruction accesses a protected memory location under a wrong assumption. Static taint analysis tools may avoid this issue. Yet, static analyses often overapproximate results and can still be incomplete in the tracking of exact data types, leading to error-proneness in the instrumentation. Any secrecy tracking that leads to hybrid (instrumented and non-instrumented) usage of global variables endangers functional correctness of the program, as detailed in Section VIII.
@@ IV Defense Implementation Layers
As previous research indicated, ciphertext side-channel defenses can be realized at different implementation layers: The ad-hoc patches to WolfSSL [42] and the Linux kernel [21] were made to the source code, while Cipherfix [37] directly modified the compiled binaries through binary rewriting. In the following, we provide guidance on a suitable implementation layer for an automated mitigation. We accumulate knowledge gained from existing side-channel defenses, explore possible solutions and discuss practical limitations. For compiler-based approaches, we generally use terminology from the LLVM project [19], which has long become a standard tool for side-channel defense implementations. However, our conclusions also apply to similar compiler infrastructures.
@@ IV-A Binary Rewriting
In binary rewriting, the modifications are directly applied on an existing binary. This approach has the advantage of being able to harden an application specifically for deployment on a vulnerable system, e.g., an SEV-SNP protected VM, without needing recompilation. Another advantage is the ability to protect not only a single application binary, but its dependencies as well, which may be too complex to compile and deploy separately (e.g., libc). The most significant challenges of binary rewriting are the stability of the modified binary, and runtime performance [37]. While binary rewriting is generally possible for typical instrumentation tasks, ciphertext side-channel defenses require very extensive changes, pushing any binary rewriting framework to its limits [13, 8].
@@ IV-C Compiler: LLVM IR/Middle End
Summary. In order to benefit from eased adjustments to memory layouts, instrumentation needs to be on middle end level. The advantages of back end level instrumentation only become apparent if the instrumentation happens after the register allocation to be able to protect all memory writes. Yet, at the back end stage, it is already sufficiently complicated to adjust the memory layout for interleaving, making it comparable to binary instrumentation. We conclude that hardening ciphertext side-channels at the IR layer promises the best final performance; and combined with a binary-level assessment, an appropriate security assurance is provided.
@@ V Zebrafix Implementation
As a case study for evaluating the described interleaving approach, we implement a proof-of-concept, dubbed Zebrafix. With that, we showcase the versatility of interleaving as a side-channel defense. Zebrafix is implemented as a set of LLVM 16111LLVM 16 is at the time of writing the newest LLVM version supported by SVF. out-of-tree passes with SVF [30] pointer analysis, with 2,6002,6002 , 600 LOC in total. An overview of the Zebrafix toolchain can be found in Figure 5. The developer who wants to protect a certain function against ciphertext side-channel leakage marks it with a $\mathtt{clang}$ function attribute. The attribute is automatically propagated through the call tree, leading to instrumentation of all called child functions as well. Indirect calls and unsupported components like inline assembly are printed for further inspection by the developer during the call tree instrumentation which is based on the LLVM call graph. We merge all bitcode files that are generated for whole-program analysis and then instrument the merged bitcode file. A discussion about the interaction of instrumented and non-instrumented (e.g., dynamically linked) code can be found in Section VIII. Zebrafix supports the block size 16 byte with data chunks up to size 16/2=816/2=816 / 2 = 8 byte. An example of data chunks smaller than 8 byte can be found in Figure 6.
@@ V-A Compiler-Level Transformations
In addition to the LLVM IR passes, we adjust the register allocator for x86-64 targets to decrease the number of stack spills of potentially secret data during function calls or because of high register pressure inside of functions. For that, we integrate a register allocator enhancement that spills to vector registers instead of the stack, proposed by Matthias Braun [^3] We adapted https://discourse.llvm.org/t/rfc-spill2reg-selectively-replace-spills-to-stack-with-spills-to-vector-registers/59630/15.. If there are not enough vector registers available for saving values, we print a warning so that the developer can check whether the write actually contains secret-dependent data or just irrelevant data like a pointer that is written to the stack. The check is further enhanced with a binary-level assessment tool. For the sake of maintainability and extensibility of Zebrafix, we chose not to implement interleaved stack spilling, which would have required deep changes to the stack allocation.
@@ V-B Adjusting Compiler Intrinsics
Examples of memory-altering intrinsics are $\mathtt{memcpy}$ and $\mathtt{memset}$, which we replace with own implementations in Zebrafix. As the 16-byte blocks that are built with Zebrafix contain varying amounts of data (e.g., 1 byte for chars, 4 byte for short integers, 8 byte for long integers), it does not suffice to adjust the amount of bytes copied or set by a fixed factor. Instead, the correct number of elements of a specific primitive type to be copied has to be determined. In the Zebrafix implementation, we use SVF to infer the types of (source and) destination objects for the memory-altering intrinsics. With that, the data can in most cases properly be extracted such that a suitable interleaved struct for the copy or set destination can be generated. The struct is then used for inserting new instrumented writes to memory. In addition, the counter values must not only be copied, but also incremented to prevent leakage from multiple write accesses with the same counter value. Thereby, even with multiple copies of the buffer, the counters are incremented to avoid repeating ciphertexts.
@@ V-C Instrumentation Aspects for Interleaving
Figure 7: Instrumentation of a 1-byte store instruction. The value to be written is combined with a counter via register-only operations. A single instruction writes both value and counter.
@@
V-D Binary-Level Assessment
To ensure that the LLVM middle end level instrumentation is not removed by compiler passes that occur later in the compilation process, and that inserted stack spills do not leak any patterns of repeating values, we implement a binary-level assessment tool. The assessment tool is a Pintool [23] that traces all memory writes of the instrumented binary and manages a set of all written values per address. Whenever there is a repetition of written values within a set, the corresponding instruction offsets and written values are logged to a trace file so that the developer can manually check them for potential leakages in accordance with known leaking algorithms. That way, the developer can determine whether secret data would be spilled to the stack and can apply some source-level mitigations like storing the affected data as global variables instead. Afterward, the compilation and binary-level check is repeated until a non-leaky version is produced (cf. Figure 5).
@@ VI Zebrafix Evaluation
In order to categorize the presented approach of Zebrafix as a mitigation against ciphertext side-channel leakage, we assess the security and performance in comparison to the Cipherfix masking approach. We perform our evaluation on a system with an Intel 4th Gen Xeon Gold 6438Y+ CPU (Sapphire Rapids) with 512512512 GiB of DDR5 and local SSD storage. We base our evaluation on target primitives from $\mathtt{libsodium}$-1.0.20 and $\mathtt{mbedTLS}$-3.6.0.
@@ VI-A Compile Time and Memory Consumption
The average code size of instrumented binaries increases by around 636363% (from 220220220/250250250 MB to approx. 290290290/435435435 MB), where a recurring part is counter management and the rest belongs to other adjustments of instructions and global variables. The average memory usage increases by a factor of five; this is mostly due to lack of optimizations like checking the heap data types or filling each 8-byte data chunk of the interleaved structs with data. The comparatively high memory usage increase in $\mathtt{SHA512}$ is an example of such missing heap data type optimizations. More details can be found in Table I.
Table I: Overhead of compilation time, code size and memory consumption. Compilation time and code size are related to the library version containing certain instrumented primitives (baseline given in the library rows), whereas memory usage is estimated for the execution of the primitive in isolation.
@@
$\mathtt{libsodium}$ 3.43.43.4 250250250
$\mathtt{Ed25519}$ 8.18.18.1 437437437 2,7442,7442 , 744 10,33610,33610 , 336
$\mathtt{SHA512}$ 8.38.38.3 $45$,6005,6005 , 600 59,56859,56859 , 568
$\mathtt{ChaCha20}$ 7.67.67.6 389389389 2,2882,2882 , 288 7,4887,4887 , 488
$\mathtt{mbedTLS}$ 2.62.62.6 218218218
$\mathtt{Base64}$ 5.85.85.8 286286286 1,5441,5441 , 544 3,4883,4883 , 488
@@ VI-B Runtime Overhead and Security
The primitives for the evaluation were chosen according to the following criteria: $\mathtt{EdDSA}$ has been shown to be vulnerable against ciphertext side-channel attacks [21, 37]. Apart from that, we include $\mathtt{Base64}$ as it was the fastest target in prior ciphertext side-channel mitigation evaluations, and $\mathtt{SHA512}$ as another fast target for comparison. We depict the $\mathtt{Base64}$ decoding in Table II; the results for the encoding are similar. All targets include zeroing of the resp. memory when freeing resources to avoid leakage of data through memory reuse. As mentioned in Section VIII, we exclude target versions that contain inline assembly or features that the current proof-of-concept does not support, like $\mathtt{AES}$ vector instructions. In the latter case, the computation mainly happens in registers, so rare leaking spots can be subject to manual adjustments.
@@ VI-C Comparison to Cipherfix Masking Approach
Zebrafix and Cipherfix introduce memory overhead from the ciphertext side-channel protection compared to baseline execution. While Zebrafix extends each data chunk with a counter in interleaved memory writes, Cipherfix is based on masking each plaintext chunk via XORing the data and the random mask. In Cipherfix, the Base and Enhanced versions contain a same-sized secrecy and mask buffer for each data chunk. In the Cipherfix-Fast version, there is only one merged buffer for each data chunk. Thereby, the memory overhead in Zebrafix is equal to Cipherfix or even smaller if 8-byte chunks are interleaved, whereas Zebrafix introduces a larger memory overhead for smaller data types (Figure 6).
@@
Even for the fastest primitive with the weakest security level in the Cipherfix evaluation, $\mathtt{Base64}$ from $\mathtt{mbedTLS}$ and Cipherfix-Fast with $\mathtt{AES}$, Zebrafix reaches a similar overhead. In Cipherfix, only secret-dependent instructions are protected. Thus, Zebrafix’s performance impact could be reduced even further with secret tracking and other optimizations, e.g., allowing the use of vector instructions combined with interleaved memory writes.
In general, our results suggest that our proposed interleaving approach is more efficient than the data masking approach taken by Cipherfix. This is due to the fact that randomness generation introduces a significant overhead that we avoid. The remaining speedup of Zebrafix over Cipherfix can be explained with a more optimized compiler-built binary, instead of relying on binary rewriting. We conclude that implementing an interleaving-based ciphertext side-channel mitigation in the compiler middle end can enhance performance and security compared to masking-based binary-level approaches.
@@
$\mathtt{Ed25519}$ 39.139.139.1$\times$ 5.55.55.5$\times$ 8.68.68.6$\times$ 1.61.61.6$\times$
$\mathtt{SHA512}$ 21.621.621.6$\times$ 1.61.61.6$\times$ 2.42.42.4$\times$ 1.31.31.3$\times$
@@
$\mathtt{Base64}$ 4.64.64.6$\times$ 1.21.21.2$\times$ 1.61.61.6$\times$ 1.11.11.1$\times$
@@ VII Preventing Leakages from Silent Stores
With the discussion of ciphertext side-channel defenses in mind, we now explore connections to silent stores to show how these memory-centric leakages can be generalized to more holistic defenses. Per se, all attacks described in Section II are based on different mechanisms to exploit leakage. However, despite originating from different sources, all of these leakages can be considered memory-centric side-channels and thus be thwarted by principled defenses: If the data was not architecturally written to memory in a leaky way in the first place, the different leakage mechanisms are no longer effective. In this section, we outline the interconnection of ciphertext side-channel and silent store leakages and compare the effectiveness of Zebrafix with silent store mitigations.
@@
In the ciphertext side-channel leakage scenario, an attacker can observe deterministic patterns of ciphertext changes of victim VM data and thereby derive information about secret keys. The attacker model of ciphertext side-channel leakage is the strongest one to be considered. In the other memory-centric leakage scenarios, we assume that microarchitectural optimizations that are currently on the rise, like the ones discussed in Section II, will gain more importance and see implementation in common processors. The attacker is able to bring the system into a state where the best conditions for maximal leakage through microarchitectural optimizations have been established; thus we assume the worst case leakage in each channel to be possible. [^4] At the time of writing, we are not aware of systems that already feature the microarchitectural optimizations in question, apart from the ones that are linked in Section II. The assumptions on code that is instrumented are in line with the ones presented in Section III-A.